OpenStudy (anonymous):

Use the quadratic formula to solve the following quadratic equation: 8x2 + 2x - 3 = 0. A. x = -7/8 or 3/8 B. x = -11/8 or 9/8 C. x = -3/4 or 1/2 D. x = -1/2 or 3/4

OpenStudy (anonymous):

It's in the form ax² + bx + c = 0, so a = 8, b = 2, c = -3 The quadratic equation is x = (-b ± √(b² - 4ac))/(2a) So you have x = (-2 ± √(2² - 4*8*(-3)))/(2*8) = (-2 ± √(4 + 96))/16 = (-2 ± √100)/16 = (-2 ± 10)/16 If it's +, you have 8/16 = 1/2 If it's -, you have -12/16 = -3/4 The answer is -3/4 or 1/2, so it's C

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

You're welcome! Do you understand it?